# FlashBias: Fast Computation of Attention with Bias

### v1

@@ Appendix C More Results in AlphaFold 3 abramson2024accurate
Analysis of pair representation bias
To give a clear illustration of the learned neural decomposition, we also plot the original pair representation bias and FlashBias approximated bias in Figure 6. From the visualization, we find that neural decomposition can give a relatively accurate estimation of the bias matrix, which performs well in capturing the “texture” of the original bias matrix. In addition, it is also observed that neural decomposition is not completely perfect in reconstructing the diagonal weights. Despite this deficiency, FlashBias still maintains the original accuracy of AlphaFold3, as presented in Table 4.4 of the main text. This may benefit from the dot-product self-attention mechanism and residual connection, which can give a robust and dominating weight for self-relation modeling.
Figure 6: Comparison between neural decomposed factor tensors’ multiplication and original bias. Here we visualize the bias of 7r6r (245 residues) and 7pzb (600 residues) in the first layer of Pairformer, which contains four heads. Their concrete 3D folding structure is plotted in Table 4.4. We also mark the rank value that can maintain 99% energy of the original bias in the above figure.
Appendix D More Results in Swin Transformer V2 liu2022swin
Figure 7: Maximum rank value of the split two subsets across 24 layers in SwinV2-B.
@@ Appendix D More Results in Swin Transformer V2 liu2022swin
As described in Section 4.3, we split the bias matrix of every layer into two subsets according to the rank value that can maintain 95% energy. Here, each layer contains a bias matrix with shape #$\#\operatorname{heads}\times 576\times 576$ and each subset is of shape #$\frac{\#\operatorname{heads}}{2}\times 576\times 576$. As shown in Figure 7, the maximum rank value of the low-rank subset is smaller than 100 in 23 layers (except the third layer). Based on these statistical data, we set $R$ in FlashBias as [32, 32, 280, 40, 48, 40, 40, 88, 80, 32, 64, 32, 32, 32, 32, 32, 32, 88, 32, 32, 32, 32, 32, 32] for the low-rank subset in each layer. Especially, we set $R\equiv 0\pmod{8}$ following FlashAttention v2 dao2023flashattention to facilitate block-wise reading.
@@
In Figure 5 of the main text, we visualize the original bias matrix at the 15th layer with 16 heads. To deliver an intuitive understanding, we also take this layer’s bias as an example and compare it with the SVD decomposed factor tensors’ multiplication. Since we only adopt SVD decomposition on half heads for speedup, we only plot 8 heads here. As shown in Figure 8, SVD decomposed factor tensors can accurately reconstruct the original bias matrix.
@@
Appendix E More Results in Transformer PDE Solver
@@
Appendix F Generalization for Diverse Biases
@@
where $\mathbf{x}_{\mathbf{q},i},\mathbf{x}_{\mathbf{k},j}$ denotes the spatial positions of the $i$-th and $j$-th objects. We train $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ based on randomly sampled points from $[0,1]\times[0,1]$ in the 2D space. Since the bias is inversely proportional to spatial distance, we further add 0.01 to the diagonal bias for numerical stability.
@@
where $\mathbf{x}_{\mathbf{q},i},\mathbf{x}_{\mathbf{k},j}\in\mathbb{R}^{2}$ records the latitude and longitude of the $i$-th and $j$-th position respectively. Similarly, we also train $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ based on randomly sampled points in $[-\pi,\pi]\times[0,2\pi]$.
For the biases mentioned above, we set $R=32$ and $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ as three linear layers with in-between tanh activation functions. Then we optimize these model parameters with the Adam DBLP:journals/corr/KingmaB14 optimizer for 10000 iterations, which will take less than 30 seconds on an A100 GPU. As presented in Figure 9, neural decomposition performs very well in these two biases, especially for the spherical bias. As for gravity bias, since the numerical instability of the inverse proportion function, it is more difficult for optimization, while our method still captures the locality of the bias matrix.
@@
Appendix G Implementation Details
As a supplement to the main text, we include more details of implementation here. All the experiments are conducted based on PyTorch 2.5.0 Paszke2019PyTorchAI and on a single A100 GPU with 144 CPU cores. All efficiency metrics are averaged from 1000 iterations. For example, some metrics are recorded as s/100iters or s/10iters, where we divide the running time of 1000 iterations by 10 or 100, respectively. In our experiments, all the algorithms’ efficiency performance is quite stable.
@@ Appendix G Implementation Details
This part of the experiments strictly follows the official code in Swin Transformer V2 liu2022swin . As illustrated in Figure 7, we only apply FlashBias to the computation of half heads in each layer (except the third layer). To further reduce the memory addressing cost, we also permute the linear layer parameters to move the speedup heads to the first half dimension, including the linear layer for generating queries, keys, values and the linear layer for final output. This operation can also be finished offline and directly applied to the inference phase.
@@
In AlphaFold 3 abramson2024accurate , we adopt the neural decomposition version. This version involves the training of newly added neural layers, whose finetuning configuration is listed in Table 10. Here we find that optimizing the factor functions for 2000 iterations can obtain a nearly converged version. This may be because the factor functions are token-wise, which means at every iteration, the model will receive $N$ (sequence length) samples for training. To sum up, these layers have been optimized with around 768000 samples after 2000 iterations. Also, to reduce the cumulative error, we only apply FlashBias to the first 16 Pairformer blocks.
@@
Output $\widehat{\phi}_{\mathbf{q},\theta_{1}}(\mathbf{x}_{\mathbf{q}})$, $\widehat{\phi}_{\mathbf{q},\theta_{1}}(\mathbf{x}_{\mathbf{k}})$ Decomposed factor tensors with shape of $N\times\#\operatorname{heads}\times 64$ (
#$\#\operatorname{heads}=4$
Model $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ Input Dim: 577; Hidden Dim: 256; Output Dim: 256 ($=4\times 64$)
@@
Overall steps: 2000 iterations

### v3

@@ Appendix E More Results in AlphaFold 3 abramson2024accurate
Appendix F More Results in Swin Transformer V2 liu2022swin
Figure 10: Maximum and mean rank value to remain 95% energy of bias matrices across different heads of 24 layers in SwinV2-B.
@@ Appendix F More Results in Swin Transformer V2 liu2022swin
We visualize the rank of bias matrices that can remain 95% energy across 24 layers. As presented in Figure 10, the bias of later layers are of lower rank. To reduce the accumulation error, we only apply FlashBias to the last 8 layers as described in Section 4.3. Specifically, for the last 8 layers, we set $R$ in FlashBias as 16. In fact, applying FlashBias to more layers will bring more significant speedup, while to maintain the model accuracy, we only apply FlashBias to the last 8 layers in this work.
@@
In Figure 6 of the main text, we visualize the original bias matrix at the 20th layer with 16 heads. To deliver an intuitive understanding, we also take this layer’s bias as an example and compare it with the SVD decomposed factor tensors’ multiplication. As shown in Figure 11, SVD decomposed factor tensors can accurately reconstruct the original bias.
@@
Appendix G More Results in Transformer PDE Solver
@@
Appendix H Generalization for Diverse Biases
@@
where $\mathbf{x}_{\mathbf{q},i},\mathbf{x}_{\mathbf{k},j}$ denotes the spatial positions of the $i$-th and $j$-th objects. We train $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ based on randomly sampled points from $0,1]\times[0,1]$ in the 2D space. Since the bias is inversely proportional to spatial distance, we further add 0.01 to the diagonal bias for numerical stability.
@@
where $\mathbf{x}_{\mathbf{q},i},\mathbf{x}_{\mathbf{k},j}\in\mathbb{R}^{2}$ records the latitude and longitude of the $i$-th and $j$-th position respectively. Similarly, we also train $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ based on randomly sampled points in $\pi,\pi]\times[0,2\pi]$.
For the biases mentioned above, we set $R=32$ and $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ as three linear layers with in-between tanh activation functions. Then we optimize these model parameters with the Adam DBLP:journals/corr/KingmaB14 optimizer for 10,000 iterations, which will take less than 30 seconds on an A100 GPU. As presented in Figure 12, neural decomposition performs very well in these two biases, especially for the spherical bias. As for gravity bias, since the numerical instability of the inverse proportion function, it is more difficult for optimization, while our method still captures the locality of the bias matrix.
@@
Appendix I Implementation Details
As a supplement to the main text, we include more details of implementation here. All the experiments are conducted based on PyTorch 2.5.0 Paszke2019PyTorchAI and Triton 3.0.0 triton on a single A100 GPU with 144 CPU cores. All efficiency metrics are averaged from 1,000 iterations. For example, some metrics are recorded as s/100iters or s/10iters, where we divide the running time of 1,000 iterations by 10 or 100, respectively. In our experiments, all the algorithms’ efficiency performance is quite stable.
@@ Appendix I Implementation Details
This part of the experiments strictly follows the official code in Swin Transformer V2 liu2022swin . As illustrated in Figure 10, we only apply FlashBias to the last 8 layers. It is also possible to apply FlashBias to part of low-rank heads in each layer. However, this will require some preprocessing steps, such as rearranging the heads and model parameters to ensure that all low-rank heads are memory continuous. Considering the implementation simplification, we eventually adopt the current design, which only speeds up the final layers.
@@
In AlphaFold 3 abramson2024accurate , we adopt the neural decomposition version. This version involves the training of newly added neural layers, whose finetuning configuration is listed in Table 12. Here we find that optimizing the factor functions for 10,000 iterations can obtain a nearly converged version. This may be because the factor functions are token-wise, which means at every iteration, the model will receive $N$ (sequence length) samples for training. To sum up, these layers have been optimized with around 3,840,000 “samples” after 10,000 iterations. Also, to reduce the cumulative error, we only apply FlashBias to the first 16 Pairformer blocks.
@@
Output $\widehat{\phi}_{\mathbf{q},\theta_{1}}(\mathbf{x}_{\mathbf{q}})$, $\widehat{\phi}_{\mathbf{q},\theta_{1}}(\mathbf{x}_{\mathbf{k}})$ Decomposed factor tensors with shape of $N\times\#\operatorname{heads}\times 96$ (
$\operatorname{heads}=4$
Model $\widehat{\phi}_{\mathbf{q},\theta_{1}},\widehat{\phi}_{\mathbf{k},\theta_{2}}$ Input Dim: 577; Hidden Dim: 256; Output Dim: 384 ($4\times 96$)
@@
Overall steps: 10,000 iterations
@@
Appendix J Extension to Multiplicative Bias
In this section, we will discuss the extension to multiplicative bias in the following formalization:
$\mathbf{o}=\operatorname{softmax}(\frac{\mathbf{q}\mathbf{k}^{\top}}{\sqrt{C}}\odot\mathbf{b})\mathbf{v},$ (15)
where $\mathbf{b}\in\mathbb{R}^{N\times N}$ represents the multiplicative bias and $C$ denotes the number of hidden channels. $\odot$ denotes the Hadamard product, namely the element-wise multiplication.
Rotary position embedding (RoPE) su2024roformer
As mentioned in Section 2.1, RoPE is a special multiplicative bias. Specifically, its pre-softmax attention weight is defined as follows:
$a_{mn}=\operatorname{Re}[\sum_{i=0}^{C/2-1}\bar{\mathbf{q}}_{i}\bar{\mathbf{k}}_{i}e^{i(m-n)\theta_{i}}],$ (16)
where $a_{mn}$ represents the pre-softmax weight between $m$-th query and $n$-th key. $\bar{\mathbf{q}}_{i}\in\mathbb{R}^{1\times 2}$ represents the $2i$ and $2i+1$-th element of the $m$-th query representation. By comparing the above Eq. (15) and Eq. (16), we can obtain the following observations:
(i) RoPE does not follow the common formalization of multiplicative bias. As formalized in Eq. (16), RoPE does not directly multiply a bias weight to each attention value, which also involves a detailed reweighing along the channel dimension. Thus, we prefer to consider RoPE as a unique technique, not a “general formalization” for multiplicative bias.
(ii) RoPE is already FlashAttention-friendly and does not need further modifications to fit the computation of FlashAttention. According to Eq. (34) in its official paper su2024roformer , RoPE can be accomplished by multiplying the rotary tensors with $\mathbf{q}$ and $\mathbf{k}$ before computing $\mathbf{q}\mathbf{k}^{\top}$. With this design, it does not need to load an additional $N\times N$ bias matrix; thereby, it can be seamlessly integrated with FlashAttention. Thus, we do not think it needs further modifications to fit the computation flow of FlashAttention. That is also why we do not experiment with RoPE in this paper.
General multiplicative bias
Here, we will demonstrate that the low-rank decomposition proposed in FlashBias can be extended to multiplicative bias defined in Eq. (15).
Specifically, suppose that $\mathbf{b}\in\mathbb{R}^{N\times N}$ can be decomposed as the multiplication of two rank-$R$ tensors: $\phi_{\mathbf{q}}$ and $\phi_{\mathbf{k}}\in\mathbb{R}^{N\times R}.$ We can rewrite the calculation of attention with multiplicative bias as follows:
$\begin{split}&\mathbf{o}=\operatorname{softmax}(\frac{\mathbf{q}\mathbf{k}^{\top}}{\sqrt{C}}\odot\mathbf{b})\mathbf{v}=\operatorname{softmax}(\frac{\mathbf{q}^{\prime}{\mathbf{k}^{\prime}}^{\top}}{\sqrt{C}})\mathbf{v},\\ \text{where}\ &\mathbf{q}^{\prime}=[\mathbf{q}\odot\phi_{\mathbf{q},1},\cdots,\mathbf{q}\odot\phi_{\mathbf{q},R}]\in\mathbb{R}^{N\times CR},\ \mathbf{k}^{\prime}=[\mathbf{k}\odot\phi_{\mathbf{k},1},\cdots,\mathbf{k}\odot\phi_{\mathbf{k},R}]\in\mathbb{R}^{N\times CR}.\end{split}$ (17)
Here $\phi_{\mathbf{q},i},\phi_{\mathbf{k},i}\in\mathbb{R}^{N\times 1}$ represents the $i$-th channel of the decomposed factor tensors $\phi_{\mathbf{q}}$ and $\phi_{\mathbf{k}}$, whose channel dimension will be broadcasted $C$ times before multiply to $\mathbf{q}$ or $\mathbf{k}$. Operation $\cdots$$\cdots$] denotes the concatenation along the channel dimension. Besides, the above computation also requires repeating the original $\mathbf{q}$ and $\mathbf{k}$ along the channel dimension for $R$ times and multiplying the factor tensors for each repeat. It is easy to verify the correctness of the above reformalization.
Example J.1.
Given the multiplicative bias $\mathbf{b}_{ij}=\cos(i-j)$, which can also be viewed as a simplified version of RoPE (Eq. (16)), $\mathbf{b}$ can be decomposed into two factor tensors with $R=2$:
$\begin{split}\phi_{\mathbf{q}}(\mathbf{x}_{\mathbf{q},i})=[\cos(i),\sin(i)]\in\mathbb{R}^{1\times 2},\ \ \phi_{\mathbf{k}}(\mathbf{x}_{\mathbf{k},j})=[\cos(j),\sin(j)]\in\mathbb{R}^{1\times 2}.\end{split}$ (18)
Based on this decomposition, we can use the above Eq. (17) for speedup, where the whole $N\times N$ bias matrix is transformed into the reweighted repeat of $\mathbf{q}$ and $\mathbf{k}$.
Corollary J.2 (Efficiency analysis).
FlashBias’s extension to attention with multiplicative bias (Eq. (15)) can reduce the storage complexity of vanilla FlashAttention when $R\leq\sqrt{\frac{{S}}{C^{2}}+1}$.
Proof.
According to dao2023flashattention , the HBM access of FlashAttention with bias is $\Theta(\frac{NMC^{2}}{S}+NM)$, it is easy to drive that the HBM access of Eq. (17) is $\Theta(\frac{NMC^{2}R^{2}}{S})$. Thus, when $R\leq\sqrt{\frac{{S}}{C^{2}}+1}$, we have:
$\Theta(\frac{NMC^{2}R^{2}}{S})\leq\Theta(\frac{NMC^{2}(\frac{S}{C^{2}}+1)}{S})=\Theta(\frac{NMC^{2}}{S}+NM).$ (19)
Example J.3 (Design practice).
Given a regular setting $C=64$ and S=S=100KB, FlashBias can achieve acceleration for attention with multiplicative bias if there exists a decomposition with $R\leq 27$.
Since the multiplicative bias is not as common as the additive bias and RoPE has already been well implemented, we would like to leave more explorations of multiplicative bias as our future work.
Appendix K Limitations
As mentioned in the main text, FlashBias’s speedup is based on the low-rank nature of attention bias. Although the low-rank biases are observed in many well-established backbones, if the bias matrix is not of low rank, FlashBias will have to trade off between efficiency and final performance.
About this limitation, one possible solution is to introduce an additional sparsity matrix as a supplement to low-rank decomposition, which is a well-known matrix approximation technique chandrasekaran2009sparse . Specifically, given the attention bias matrix $\mathbf{b}$, this method formalizes the decomposition process as a convex relaxation problem, which can be formalized as follows:
$\begin{split}\widehat{\mathbf{r}}+\widehat{\mathbf{t}}&=\arg\min_{\mathbf{r},\mathbf{t}}\|\mathbf{r}\|_{\ast}+\gamma\|\mathbf{t}\|_{1},\\ &\text{s.t.~}\mathbf{r}+\mathbf{t}=\mathbf{b},\end{split}$ (20)
where $\|\|_{\ast}$ represents the nuclear norm and $\mathbf{r}$ is expected to be low-rank and $\mathbf{s}$ is optimized to be sparse. Afterwards, the low-rank matrix $\mathbf{r}$ can be sped up through FlashBias, and we can adopt the same technique as attention mask speedup sharma2024efficiently for accelerating the sparse component $\mathbf{t}$. Further, this technique can also make up for the current shortage of FlashBias in approximating the diagonal value (Figure 7), where we can decompose the bias into the sum of a low-rank and a diagonal matrix.
Since we mainly focus on the low-rank property of attention bias in this paper and it will not affect the final performance in our experiments, we would like to leave this as our future work.
Appendix L Broader Impacts
[…]